For my finals I only have about 6 questions that I am stuck on. Any help is appreciated.
1) write quadratic equation in variable with a solution of 4 only.
2)Identify the degree of each term of the polynominal of -8x^3+2x^2+2x+4.
3)rewrite the expression with positive exponents (19xy)^-9/10
4)evaluate the polynomial for x=6 for 3x^2-2x+5
5) convert to decimal notation 8.87*10^7
6)find the variation constant and an equation of variation where y varies directly as x and y=90 when x=10
Thank you sooo much for any help!!!

Answer by stanbon(75887)

1) write quadratic equation in variable with a solution of 4 only.
Ans: f(x) = (x-4)^2 = x^2-8x+16
-------------------------------------------
2)Identify the degree of each term of the polynominal of -8x^3+2x^2+2x+4.
Ans: from left to right: 3,2,1,0
--------------------------------------------
3)rewrite the expression with positive exponents (19xy)^-9/10
Ans: [1/(19xy)]^(9/10)
----------------------------------------------
4)evaluate the polynomial for x=6 for 3x^2-2x+5
Ans: f(6) = 3*6^2-2*6+5 = 3*36-12+5 = 108-7 = 101
-----------------------------------------------
5) convert to decimal notation 8.87*10^7
Ans: 88,700,000
-----------------------------------------------
6)find the variation constant and an equation of variation where y varies directly as x and y=90 when x=10
Ans: y = k*x
90 = k*10
k = 9
-------------------------------
Cheers,
Stan H.
===================